### v1

# Phase segregation lubricates paste extrusion

Abstract
Extrusion is a critical process in the manufacturing of various organic materials, including animal feed [1], human food [2], pharmaceuticals [3], and biomass for bioenergy [4]. During extrusion, a dense particle-fluid mixture, or paste, is pushed through a narrow channel under high pressure [5, 6]. Despite its widespread use, many questions remain about how particle-fluid interactions and frictional forces at the wall influence high-pressure paste flow [7, 8], particularly whether laboratory-scale findings translate to more complex industrial settings[9]. Here, we quantify for the first time how a micrometer thin fluid lubrication layer emerges from particle-fluid interactions during the extrusion of organic biomass, controlling friction at the paste-channel interface. This lubrication layer forms through a pressure-induced phase segregation process [10, 9, 11] and is significantly thinner than the particle size [12], affecting their contact dynamics, making friction at the interface tunable. Our results reveal how the thickness of the lubrication layer is influenced by paste composition, enhancing our fundamental understanding of phase segregation during industrial extrusion processes and offering potential pathways for reducing energy consumption and product failure  [13] in the industrial material production.
preprint: APS/123-QED
The flow behavior of concentrated suspensions and pastes remains a challenge to understand, affecting processes ranging from human food production and additive manufacturing to medical and clinical applications [10, 14, 15, 16, 17, 18, 11, 19, 2]. Pastes are mixtures of particles and a small amount of liquid. During their extrusion, these two-phase mixtures of particles and liquid are pressured into a pre-shaped channel, the die, to obtain a final product. Pressures used are 100s of MPa, high enough to induce particle deformation and bonding [20]. Optimization of the material and factory performance during this process happens via paste ingredient adjustments, die shape modification and process temperature and pressure alteration  [21, 6] and is performed at a daily basis in many plants worldwide. Our fundamental microscopic understanding of the typically non-Newtonian flow behavior of pastes [7, 8] during high pressure extrusion is however very limited. Controlling the extrusion process is challenging due to the numerous tuning parameters involved [5, 6], yet currently there is no physico-chemical framework that allows us to understand their interrelatedness. Here we provide one new fundamental paste-flow mechanism that binds many extrusion parameters together. We show that many extrusion tuning parameters are fundamentally linked in their ability to tune an ultrathin liquid film that emerges during extrusion. The film is only micrometers thick, as it is located between the extrusion die and the extrudate. The frictional role of the film so determines the mechanical energy usage and temperature evolution during the extrusion process, thereby also controlling a large number of product quality parameters.
During paste extrusion, physical forces act on the solid-liquid mixture, deforming the particles and causing the solid particles and liquid fraction to undergo relative displacement. These phenomena often referred to as “migration”, “demixing” or solid-liquid phase separation. In general, such solid-liquid separation is known to result in flow blockages, product inhomogeneity and product failure [17, 22, 23, 13]. It was conjectured long ago that also a beneficial liquid-rich layer can form between the wall of the die channel and the flowing paste [10], resulting in self-lubricating paste flow. Such liquid lubrication layers are known to change sliding resistance by orders of magnitude [24, 25], particularly in the so-called “mixed”-regime where a lubrication layer starts to reduce the solid-to-solid sliding friction [26]. As the lubrication layer affects friction or paste tribology, it influences the mechanical energy use and heat generated inside the extrusion channel at the (sub)micrometer length scale. In the extrusion die, precise tuning of both energy and temperature are necessary to induce and control agglomeration of the particulate mixture through various physical and chemical transitions, such as chemical reactions, melt transitions, and sintering [20]. Control over these transitions is crucial in the formation of a high-quality end product [20, 6, 1, 27, 2]. The control over phase segregation induced lubrication is thus an essential step in paste extrusion. Our post-extrusion surface liquid evaporation measurements quantify this phase segregation by measuring the lubrication film thickness, allowing for deep insight into the flow dynamics of pastes under pressure.
Figure 1: Lubrication layer thickness increases with water use during extrusion of biomass pellets. a, During pellet manufacturing, ingredient mixtures are extruded through holes within a metal ring-die, forming cylindrically shaped pellets. The water from the steam conditioning ends up on the surface of the pellets in the high pressure extrustion due to solid-liquid phase segregation. The thickness of the lubrication layer is determined immediately after extrusion through a controlled evaporation measurement on freshly sampled pellets, recording sample mass and relative air humidity ($\mathrm{RH}$). b, Evaporation measurements show that an initial saturation period ($\mathrm{RH}>95$%, denoted as $t_{sat}$) is present, arising from the evaporation of water exposed at the pellet surface. c, The mass-detected average amount of evaporated water depends on the amount of steam added during the conditioning step: the change in water content ($\Delta\mathrm{mc}$) of $0.057$ (top), $0.033$ (middle), and $0.009$ (bottom), given in kg water per kg ingredients. The bold lines highlight the amount of water that evaporates during $t_{\mathrm{sat}}$ as indicated in (b) and hence quantifies the amount of surface water that evaporates from the pellets exterior surfaces. d, The total surface water evaporation indicates the lubrication layer thickness. The layer thickness is micrometers thick and increases as more water is added in the steam conditioning process as captured by $\Delta\mathrm{mc}$. e, Surface topology scans of dried pellets confirm thickness calculations, as they do not show striations along the extrusion direction when the lubrication layer is thicker than the intrinsic height of the asperities in the metal die (top). Decrease of the lubrication layer thickness results in clear striations on the pellets surface along the extrusion direction—running left to right (bottom). Scale bar: 50 $\mathrm{\mu m}$. These images confirm that the metal die and the pellet are largely separated by the lubrication layer, schematically represented in (f).
The lubrication layer thickness was suggested to be only micrometers thick [10, 7, 8, 2], a layer too thin to measure directly in any realistic extrusion process. It has thus so far been impossible to understand how this lubrication layer emerges or depends on ingredients and process parameters. Various experimental methods have been developed to analyze phase segregation in extrusion processes at laboratory scale [10, 5, 28, 29, 30, 31, 32]. These studies revealed that flow behaviour during extrusion is highly dependent on ingredient and process variables, including solid-liquid mixing ratio, particle size, deformability of particles, flow rate, pressure, and viscosity [15, 5, 33], yet stopped short of directly showing the existence of such a lubrication layer, let alone quantify it. Our surface liquid evaporation measurements [34], directly reveals (1) that this lubrication layer exists, (2) that it is indeed only several micrometers thick, and (3) we provide evidence that the thickness is set by process parameters. We (4) confirm that lubrication tuning indeed results in frictional changes by observing energy savings and temperature modulation. We demonstrate our results via pilot-scale pelleting plant trials, with realistic organic ingredient streams as typically used in biofuel and animal feed production. Our results can be translated to similar unit operations used in ceramics, foods, and pharmaceuticals production.
I Water induced lubrication in pelleting processes
Pellets are formed in a unit operation known as the pellet extrusion process (Fig. 1a&f), where particle mixtures, often from multiple ingredients, are pressed by a roller through cylindrically shaped holes in a metal ring-die. The pellet extrusion process transforms the loose and low-density particle mixture into densely packed solid pellets, which are subsequently cooled and dispatched to the end-user. Unlike hard inorganic particle pastes used in e.g. concrete extrusion, organic plant-based particles utilized in food and feedstuff production exhibit hygroscopic properties, are capable of binding or releasing water, and undergo deformation and physico-chemical transitions while being processed. Consequently, the interaction mechanism and physical forces at the boundary layer between the metal extrusion die and the food/feed paste [2] are expected to differ from those of hard particles which for example are subject to excluded volume effects at the die-wall[7, 5, 8, 35].
Prior to pelletization, conditioning uniformly mixes the ingredient particles and steam. The steam condenses upon cooling, transferring its latent heat into the particle mix and forms water [1, 36, 37]. Consequently, the steam conditioning process sets the temperature and amount of water present in the paste before it enters the extrusion die. It is known that increasing the amount of water in the mixture reduces the mechanical energy required per kilogram of produced pellets [37]. While the water-induced lubrication in pellet extrusion is evident, the underlying mechanism remains unclear. The observed layer is too thin to be related to a particle migration phenomena, although such effects can also significantly change effective paste viscosity [38, 39]. Hitherto, one main assumption is that the paste during pellet extrusion internally retains its uniform moisture distribution from right after the mixing/conditioning step [40, 41]. We show that this is not the case: the lubricating effects of added water is caused by the formation of a thin lubricating film that partially fills the pores between the metal die and the sliding pellet (Fig. 1f), reducing the real contact area between those respective surfaces, as expected in the mixed-lubrication regime [42]. Our perspective also explains the observed reduction in equipment wear with added water [37].
Figure 2: Experimental and numerical drying results support the hypothesis of a lubrication layer. Two types of pellets underwent drying using our model drying setup (solid lines), complemented by results obtained through numerical modeling (dashed lines). The results are presented as the pellet moisture ratio, the normalized mass decay (unitless) over $900$ seconds of drying (see Eq. 6). In yellow: results corresponding to the experimental drying of pellets with a uniform moisture content, achieved through vapor sorption (stored for 1 week at over $95\%$ RH), and numerical simulation considering an initially uniform moisture distribution. In blue: results corresponding to the experimental drying of fresh pellets immediately after extrusion and numerical simulation for a pellet with a high moisture content localized at its surface, as schematically depicted. The initial drying rate is strongly affected by the presence of surface water, resulting in a quick drying profile, in contrast to pellets with a uniform moisture distribution, for which the rate of drying is determined by the transport by diffusion. Additional results are shown in Fig. A.1.
II Formation of lubrication layers during paste extrusion
We provide evidence for the existence of the lubrication layer through a pilot-scale pelleting trial. Approximately 18 tons of pellets were produced while systematically adjusting the conditioning settings, as described in the Methods section. The fresh pellets were sampled immediately after extrusion and placed in a custom designed sample container permitting drying of the product (Fig. 1a). The drying method involves exposing a sample container of pellets to a continuous, known flux of fresh air and recording the inlet and outlet air temperature and relative humidity (RH), along with the sample mass [34]. In 14 out of 15 drying experiments, a saturation period was observed ($t_{\mathrm{sat}}$), during which the passing air reached saturation (Fig. 1b). Over this period, up to 40% of the water introduced by conditioning was lost through evaporation (Fig. 1c), corresponding to the moisture “flash-off” commonly observed in pelleting factories [1]. Our modeling demonstrates that the saturation of the drying air and the initial rate of drying, observed during the drying of freshly produced pellets, can only occur if the evaporating water comes directly from the pellet surface. This is in contrast to pellets with an elevated but uniform moisture profile, which can be generated through vapor sorption (see Fig. 2; we refer to Fig. A.1 b&c and SI Section: A.2 for a detailed description of the numerical model and the results).
Evaporation of the lubrication layer thus causes the initial saturation of drying air (indicated in Fig. 1b). The amount of water evaporating during this period can be quantified as we know the total surface area of the pellets in the drying experiment from the drying mass, density and shape. We so obtain the average water layer thickness, $\lambda$, on the pellets surfaces (Fig. 1d). The thickness of the lubrication layer increases from a few micrometers, similar to the height of surface asperities on the metal die (Fig. 1e) in which case we expect frictional interactions that are also known as Stribeck boundary lubrication [24]. As we increase water content of the paste ($\Delta\mathrm{mc}$) by increasing the steam flow in the conditioning process, the layer thickness increases to several tens of micrometers. Compared to the smallest individual particles produced by the milling procedure, this layer remains very thin [12] (Fig. 1d). The water layer thus only partially separates the pellet from the die, reducing the area of direct contact between the pellet and die surface and so leads to less damage to the particles embedded in the pellet surface (Fig. 1e&f and SI section: A.4) and to less wear in the die. As the lubrication layer grows, we so expect to enter the “mixed regime” lubrication in the Stribeck perspective, and a concomitant reduction in friction. The presence of the lubrication layer also introduces plug flow in the compressed biomass as there is no visible internal deformation, uniform density and negligible porosity (see supportive X-ray tomography evidence described in SI Section: A.3). These collective findings suggest that the lubrication layer in organic feedstuffs may not arise from factors like particle size distribution and the excluded volume effect, as observed in hard particle pastes. Instead, it appears to be influenced by the roughness of the die wall. In this scenario, water, with a much lower viscosity compared to the feedstuff, flows into the die’s cavities and reduces the shearing of the feedstuff at the paste-die-wall interface by occupying part of the cavity volume, thus making it inaccessible for the feed particles. Consequently, the real contact area between the paste and metal die reduces, and the energy use decreases, as shearing of the water layer is much more favorable than shearing of the feedstuff material along the rough metal wall. This hypothesis aligns with industrial findings indicating that new pellet dies require more energy during pellet extrusion than used dies [43]. We attribute this energy use reduction over time to a decrease in the paste-die-wall contact area as the die’s largest asperities smooth out through wear. By smoothing of the surface, the water film can cover a larger area (the ratio of $\lambda$ over the wall roughness increases), further reducing the real contact area between the paste and die [25].
The ability to measure and monitor the lubrication layer thickness enables us to decouple process dependent parameters such as steam use, die-hole dimensions and production rate (the latter, together with die-geometry specifications, determines residence time) from ingredient mixture dependent parameters, such as its chemical composition, particle size, particle size distribution and thermomechanical properties [44]. We identify two important consequences: we can now identify the source of mechanical energy use of pellet extrusion—the friction at the paste-die-wall interface—and consequently reduce it. Second, by tuning the lubrication layer, we can adjust the amount of heat added during extrusion, giving control over product treatment and quality.
III The tribological role of the lubrication layer during pellet extrusion
During pellet extrusion, the sliding friction between the pellet and the metal die contributes to the energy use of the pellet press. This energy use is proportional to the extrusion pressure and increases exponentially with the compression ratio—the die-length over the die-diameter ratio—of the extrusion die $L/D$ and with the apparent coefficient of friction between the die and the paste, $\mu$ [45, 46, 47] (see SI section A.4). The dry solids friction is determined by the paste composition and its interaction with the die’s metal surface [48, 49]. In the case of animal feed, this composition is set by the raw ingredient mixture. This dry coefficient of friction can not be actively manipulated during pellet extrusion, yet the formation of a thin lubrication layer can contribute to a smooth transition from the “boundary lubrication regime” into the so-called “mixed regime” [25], significantly reducing the coefficient of friction (see SI section A.4). The decrease in the apparent coefficient of friction, $\mu$, has a direct impact on the specific mechanical energy use of the pellet press—the amount of mechanical energy required for the production of a kilogram of pellets (Fig. 3b). Hence, adjusting the lubrication layer through steam addition enables modification of the friction coefficient within a running production line, without altering the ingredient mixture composition or press/die configuration. As a consequence of the lubrication, frictional heat changes the temperature gain in pellets over the extrusion die, here called $\Delta T_{\rm die}$. We show this effect in Fig. 3c, which affects final product properties (such as “durability” as detailed in SI section A.5). An increase in extrusion rate also suppresses friction and hence SME, in accordance to Stribeck expectations [25] (see SI Fig. A.8) Temperature fluctuations can change the physico-chemical properties of the organic ingredient streams, which may result in gelatinization and melting of starches and denaturation of proteins, which subsequently affect the nutritional quality and availability during consumption [36]. Measurement and subsequently tuning of the lubrication layer is thus essential for further optimization of pelleting processes as it enables a more precise control over the mechanical energy use and product temperature compared to the current set of operating variables.
Figure 3: Lubrication layer affects the extrusion energy balance. a, Lubrication reduces the apparent die wall friction $\mu$ and, therefore, (b) reduces the net specific mechanical energy $(\mathrm{SME_{net}})$ use with increasing lubrication layer thickness $\lambda$. The reduction in mechanical energy use in particular (c) reduces die heating through friction and consequently affects the change in pellet temperature $\Delta T_{\rm die}$ over the die—the difference between the pellet’s temperature after extrusion and the ingredients temperature before extrusion, as well as final pellet temperature. The horizontal error-bars in (b) and (c) correspond to the minimum and maximum values for $\lambda$ as represented in Fig. 1d. The vertical error bars represent minimum-maximum range of at least three samples during experimental runs (see Methods section).
IV Conclusions
We confirmed the existence and thickness of a micrometer-thick lubrication layer that forms via particle-fluid phase segregation during the high pressure extrusion of pastes. We show that this lubrication layer directly affects the tribological properties of the paste flow. We used realistic, plant-based materials inside a an industrial-scale pellet extruder to evidence how the lubrication layer interlinks process variables during industrial extrusion. We so unravel the fundamental mechanisms behind a decades old observation on the lubricating effect of water addition during pellet production [37]. Our work reveals that the friction within the die is directly influenced by the thickness of the lubrication layer, providing insights for future monitoring of product quality, shape, and performance in food, feed, and pharmaceutical extrusion processes and adjacent extrusion disciplines which process inorganic ingredient streams like ceramics. Our work provides a new framework to understand and use the role of other process parameters such as die geometry, particle size distribution, and ingredient composition. Moreover, continuous monitoring of the extrusion process offers a novel approach for process control, with immediate practical implications for optimizing industrial production on a global scale. Our findings can lead to significant energy reductions, contribute to a more sustainable extrusion process and enhanced quality control in the production of a wide range of extrusion-based materials.
Acknowledgements.
We gratefully acknowledge discussions with the “Pelleting in Circular Agriculture” consortium members. We thank the Wageningen Electron Microscopy Center (WEMC) of Wageningen University & Research for access to facilities We thank the Van der Waals-Zeeman Institute for providing access to the Keyence VK-X1000. We thank Zorheh Farmani and Jordi Rijpert for their help during the experimental investigation. We thank Remco Fokkink for helping with the development of software to read out the relevant sensors during the drying experiments. The authors thank Synchrotron SOLEIL for the ANATOMIX beamline time (proposal n° 20220272). ANATOMIX is an Equipment of Excellence (EQUIPEX) funded by the Investments for the Future program of the French National Research Agency (ANR), project NanoimagesX, grant no. ANR-11-EQPX-0031.
Declarations
  • Funding: this study was financially supported by The VICTAM Foundation, Agrifirm NWE B.V., DSM, Elanco Animal Health, Pelleting Technology Netherlands, Phileo S.I. Lesaffre, Topsector Agri & Food, Zetadec, and Wageningen University & Research, as partners in the project “Pelleting in Circular Agriculture” (project number: LWV1965)
  • Conflict of interest/Competing interests: MT owns a company that provides consultancy services for the animal feed industry.
  • Availability of data and materials: all processed data is available in the main article and supplementary materials; raw data will be provided upon reasonable request.
  • Authors’ contributions:
Richard Benders: Conceptualization, Methodology, Investigation, Formal analysis, Data Curation, Visualization, Writing - original draft Menno Thomas: Conceptualization, Methodology, Writing - review & editing, Supervision, Project administration. Thomas Bastiaansen: Investigation, Writing - review & editing. Raoul Fix: Conceptualization, Methodology, Investigation, Writing - review & editing. Mario Scheel: Conceptualization, Methodology, Investigation, Writing - review & editing. Guido Bosch: Conceptualization, Writing - review & editing, Project administration Sonja de Vries: Conceptualization, Writing - review & editing Jasper van der Gucht: Conceptualization, Writing - review & editing, Formal analysis, Supervision Joshua Dijksman: Conceptualization, Methodology, Writing - original draft, Writing - review & editing, Supervision.
Appendix A Extended Data
Figure A.1: Drying experiments differentiate between fresh (quick drying) and rehydrated (slow drying) pellets, complemented by numerical results. The effect of radial phase segregation on pellet drying behavior was evaluated both experimentally and numerically. We compared the drying behavior of rehydrated pellets with a uniform moisture distribution and freshly produced pellets with increased surface wetness. a, A schematic representation illustrates the initial radial moisture distribution used as the initial condition in numerical modeling of the drying process for freshly produced pellets. The moisture ratio (unitless) represents the normalized mass decay over $900$ seconds of drying (see Eq. 6). Numerical simulations provide a qualitative comparison of the effect of radial phase segregation on drying behavior, shown in both numerical simulations (b) and experimental drying data (c) for fresh (blue lines) and rehydrated pellets (yellow lines) using the moisture distributions as the initial condition (a). The difference in drying behavior is caused by radial phase segregation, specifically the formation of a lubrication layer during pellet extrusion.
We examined the pellet drying behavior, internal and external pellet microstructure, and surface topology of pellets produced under systematically varied process conditions, adjusting only the amount of added steam to the ingredient mixture when formulating the paste, thereby changing its initial water content prior to compaction, $\Delta\mathrm{mc}$. Our objective was to determine whether phase segregation occurs during pellet extrusion. In the processing of pastes and other dense particulate suspensions, such phase segregation processes are also referred to as “spatial demixing,” “liquid migration,” “phase redistribution,” “drainage,” or “self-filtration” [11, 33]. According to our hypothesis, a thin lubricating film forms between the pellet and the metal die due to phase segregation (Fig. 1a), which fills the asperities between the two sliding surfaces and reducing the frictional contact area. As a result of the radial phase migration, the moisture distribution across the pellets becomes non-uniform, with the exterior surface of the pellet facing the die wall experiencing increased moisture levels during extrusion compared to the pellet’s interior volume [22]. Through the combined circumstantial evidence presented below, we conclude that such phase migration occurs during pellet extrusion.
While the quantification of the lubrication layer thickness is detailed in the main text, here we additionally demonstrate how adding water in the form of condensed steam reduces the apparent coefficient of friction within the mixed lubrication regime. We also discuss how tuning the lubrication layer can be performed to produce high-quality pellets with improved energy efficiency, monitoring physical pellet characteristics, and measuring the energy use within the pelleting mill. This involves balancing out the thermal energy required to produce steam and the mechanical energy required during pellet extrusion.
A.1 Pellet production
The extrusion of pellets through individual die-holes occurs semi-continuously. The ring-die rotates at a constant velocity; however, the rollers, fixed in position with respect to the rotating die, periodically pass the openings of the die-holes as the die rotates, causing compression of the ingredient mixture and subsequent extrusion of the material through the die-hole. The compression frequency, representing the number of compression steps per second, and the shear rate of the extrusion process are determined by the mass-flux of the ingredient mixture into the die, the rotational velocity of the die, and the number of rollers. The length of the pellets is controlled by knives fixed at a specified distance outside from the rotating die. Pellets shorter than this specified distance pass underneath the knives until they reach the required length, at which point they are sheared off upon collision with the knives.
A.2 Water distribution inside pellets
The concept of phase segregation during extrusion of multi-phase ingredient mixtures is not unique to pellet extrusion and occurs under a wide range of process conditions. Ranging from low viscosity soft matter suspensions, e.g. in food [50] or human health [11] applications, as well as during the extrusion of highly viscous flows, like rock melts [51]. As a result of the physical forces, an uneven distribution of the liquid fraction occurs across the product during or after extrusion. However, direct analysis of liquid distribution inside extrusion channels using non-invasive techniques like NMR and MRI at industrial process conditions is currently not possible.
Extrusion processes occur within metal dies typically made of hardened stainless steel, which can withstand high processing temperatures and pressures. This prevents the use of non-invasive experimental techniques [29, 31], along with the significant amount of dust produced in pelleting facilities which interferes with the non-invasive experimental techniques.
To address these limitations, we conducted a numerical evaluation of how radial redistribution of water over the pellet influences drying characteristics immediately after extrusion. The numerical results were qualitatively compared with experimental drying data obtained in the pilot plant.
In the numerical simulations, the initial moisture distribution along the radial coordinate $r$ was governed by a half-Gaussian function $M(r)$—a Gaussian centered at the pellet-air interface ($r_{p}$).
$M(r)=M_{\rm surf}\exp\left[-\frac{1}{2}\left(\frac{r-r_{p}}{\sigma}\right)^{2}\right]$ (1)
[…]
At the pellet’s surface, water evaporation is driven by the vapor pressure difference between the water exposed on the pellet’s surface and the vapor pressure of the air passing through the bed of pellets at a given temperature [52]. If a water layer is present on the pellet’s surface, the maximum theoretical evaporation rate from this surface can be determined using, for example, the Hertz-Knudsen-Schrage equation based on classical kinetic theory or alternatively, equations based on statistical rate theory [53]. However, during the drying experiments, we initially observe the saturation of the air passing through the bed of pellets (see Fig. 1b), leading us to conclude that the evaporation rate from the pellet’s surface is not limited by the maximum theoretical evaporation rate but is limited by the vapor capacity of the influx of fresh air. In other words, the initial evaporation of the surface water—the evaporation of the lubrication layer—occurs at such high rates that the passing air reaches its saturation point as long as sufficient water is available on the pellet’s surface. When the water layer has evaporated, the rate of evaporation reduces, as indicated by the change in slope in Fig. 3c.
Consequently, for the numerical model, mass transfer coefficient ($k_{c}$) and effective diffusion coefficient ($D_{\rm eff}$) were selected in a way that the transport rate at which water moves within the pellet (diffusion of water) is much lower than the rate of evaporation (mass transfer rate representing the change of liquid water on the surface into water vapor within the passing air). This is associated with a mass Biot number ($\mathrm{Bi_{m}}$) much larger than $1$ (see Appendix C.1). With a mass Biot number much larger than $1$, we ensure that the rate of evaporation at the surface does not limit the drying process, as the maximum evaporation rate is lower than the vapor capacity of the passing fresh air, as observed in the experimental data (see Fig. 1b).
Here, we show that the experimental drying behavior complements the expected numerical behavior when significant phase migration occurs during pellet extrusion.
The results obtained through numerical modelling (Fig. A.1b) show how a shift in water distribution, ranging from uniform (yellow) to an increased exposure of surface water (blue), affects the expected drying behaviour. Experimentally and numerically, we were able to distinguish two different cases by qualitative comparison. First, the case in which rehydrated pellets with a uniform moisture content were dried in a control experiment and secondly those of fresh sampled pellets, directly after extrusion (Fig. A.1c).
In the first case, a control experiment was performed in which pellets were rehydrated through hygroscopic vapor sorption. The moisture content throughout these pellets can be assumed near-uniform due to the long equilibration time (stored 1 week at >95% RH) [54] and consequently the drying process resembles that of a finite cylinder with surface evaporation ([55]). Here, the drying rate is controlled by diffusion through the sample (yellow lines in Fig. A.1b and Fig. A.1c). During the control experiment, no saturation of the passing air was observed. The same holds true for the simulation, during which the calculated initial rate of evaporation averaged over 10 seconds ($\approx 10^{-8}\ \mathrm{kg\ m^{-2}s^{-1}}$) was much smaller than the maximum evaporation rate based on the vapor capacity of fresh air ($\approx 10^{-4}\ \mathrm{kg\ m^{-2}s^{-1}}$) flowing through a bed of pellets and along the pellets’ surface. The latter was estimated using the vapor capacity of fresh air ($10^{-2}\ \mathrm{kg_{vapor}\ m^{-3}}$), the volumetric flow rate of fresh air into the bed ($10^{-3}\ \mathrm{m^{3}s^{-1}}$), and the total surface area of the pellets within the sample container ($10^{-1}\ \mathrm{m^{2}}$). This indicates that the evaporation rate from a bed of pellets with a uniform moisture content is much smaller than the rate required for saturation of the air. Increasing $D_{\mathrm{eff}}$ to $10^{-7}\ \mathrm{m^{2}s^{-1}}$ in the numerical drying model increases the initial evaporation rate up to $\approx 5\times 10^{-7}\ \mathrm{kg\ m^{-2}s^{-1}}$, but this still falls short of the required evaporation flux for saturation of the air, despite using a diffusivity value several decades larger than the typical reported experimental values [56, 40, 41].
In the second case, freshly sampled pellets were sampled and dried directly after extrusion. In contrast to the drying process of pellets with uniform moisture content, freshly sampled pellets exhibit a distinctly different drying behaviour (blue lines in Fig.A.1b and Fig. A.1c). The initial drying rate for fresh pellets is notably faster during both experimental and computational drying compared to the uniform initial condition driven by diffusion.
The qualitative comparison between the numerical model and the experimental drying data supports the hypothesis that phase segregation occurs during pellets extrusion. Combining the quick initial drying of the fresh pellets together with the observation of air saturation (Fig. 1 b), we conclude that water must be present in close proximity to the pellet exterior surface as a result of phase segregation during pellet extrusion.
A.3 Pellet microstructure
During pellet extrusion, the finely ground organic ingredients are bound into a cylindrical shaped pellet. Such organic ingredient streams deform easily at the typical process pressures, resulting in densification of the ingredient mixture and subsequently enhances the agglomeration process through various bond types [20]. The mechanical properties [57, 58], as well as the glass and melt transition temperatures of such organic plant derived products are strongly influenced by the local moisture content [6, 36]. Consequently, the local increase in moisture content due to the presence of a water rich lubrication layer at the pellets surface is expected to affect the surface microstructure of the pellet. Hence, we employed x-ray tomography at the synchrotron SOLEIL, utilizing the ANATOMIX beamline [59], to analyze the pellet microstructure. This analysis focused on identifying density gradients and structural changes associated with shear deformations and moisture gradients. Such gradients and structural changes can offer insights into the flow profile inside the extrusion die and the local moisture content during extrusion.
A.3.1 A uniform radial pellet density profile
At the ANATOMIX beamline, a monochromatic X-ray beam of 16.83 keV from a DCM (double crystal monochromator), at a distance of 20 cm (sample to scintillator; pixel size of 3.07 µm; exposure time 50 ms) was used to analyse the radial density profile of pellets, avoiding beam hardening artifacts. Such artifacts are typically observed in commercial x-ray tomographs, in which hardening of polychromatic x-ray beams results in a concave apparent density profile after image reconstruction. One of our hypothesis was that we expect a density gradient in a narrow region close to the pellets surface. In this region the increased moisture level could plasticise the organic material [57], making it more deformable and consequently a higher density can be obtained during the consolidation process. The use of a monochromatic beam, prevents beam hardening and enables a qualitative analysis of the radial pellet density.
Figure A.2: The pellets exhibit uniform radial density. Radial density profiles of pellets with a diameter of 6 mm, obtained through analysis of x-ray tomograms obtained at the ANATOMIX beamline, for two process conditions with (a) $\Delta\mathrm{mc}$ $0.009$ (low water content) and (b) $0.057$ (high water content) respectively. The light gray lines in (a) and (b) correspond to individual slices along the 3D volume. The black lines in (a) and (b) correspond to the average density profile over each of the slices within the stack (see Materials & Methods). The region marked by the gray area, corresponds to the area in which we observe a phase-contrast enhancement artifact. In this region the apparent density could not be determined accurately.
The apparent density near the surface is subject to phase-contrast enhancement, characterized by a negative apparent density on the exterior (outside) and a positive peak on the interior (inside) of the pellets (gray region Fig.A.2). While phase-contrast enhancements are commonly employed to enhance sensitivity, conducting a detailed study of the outer $100\mathrm{\mu m}$ region near the pellet surface—where we anticipate density changes due to high moisture contents from the lubrication layer—was not feasible due to the phase contrast enhancements. The remainder of the pellet appears to be of uniform density along the radial coordinate, at both high and low moisture levels during production (Fig.A.2).
A.3.2 Surface melt structure
Complementary scanning electron microscopy (SEM) images of the surface, show the formation of a gelatinized network of starch at the pellets surface (see Fig. A.3). Water, absorbed by the organic particles, acts a plasticizing agent and reduces the glass and melt transition temperature of starch and contributes to pellet binding [36, 6]. Hence, the starch melt is not coincidentally formed at the location where the lubrication layer forms during extrusion, in contrast to the much more limited melt formation on the interior of the pellet (see Fig. A.4). The increase in moisture content at the pellet surfaces, acts as a plasticizer and enables a localized lowering of the melt point, effectively glueing the individual particles together at the surface.
Figure A.3: Scanning Electron Microscopy images of the pellet’s surfaces (left - low water content $\Delta\mathrm{mc}=0.009$ and right - high water content $\Delta\mathrm{mc}=0.057$) demonstrate that a network of molten starch granules has formed on the pellet’s surface during pellet extrusion. At the pellet’s surface, the local moisture content has increased, lowering the melting point temperature of the starch granules to the process temperatures reached during pellet extrusion.
Figure A.4: Scanning Electron Microscopy images of the pellet’s interior (left - low water content $\Delta\mathrm{mc}=0.009$ and right - high water content $\Delta\mathrm{mc}=0.057$) demonstrate that only a limited number of starch granules on the inside of the pellet have reached their melting point temperature, while most remain in their crystalline shape. The moisture content on the interior of the pellets is expected to be lower than on the pellet’s surface due to liquid phase migration. Consequently, the starch granules on the inside have a higher melting point compared to those at the surface, and only limited melting of the starch occurs during pellet extrusion.
A.3.3 Internal pellet deformation
It has been previously proposed that the ingredient stream moves as a plug inside the die channels [60], with a uniform flow velocity through the pellet die. A lubrication layer should facilitate such plug-flow and therefore, signs of internal deformations caused by velocity gradients should be absent. A qualitative analysis of x-ray tomograms confirms this absence of internal deformations along the longitudinal axis as indications of the flow direction (Fig. A.5). Additionally, the overall porosity is low (Fig. A.5), suggesting that the ingredients are forced into place, approaching unit density during pelletization. Indeed, the estimated process pressures during pelletization [61], are of similar order, or larger than the typical compressive strength of the particles [62] and hence the particles are compressed into a tightly packed pellet with a low porosity. These observations support the hypothesis of phase segregation, wherein water in between the particles, is expelled from the ingredient matrix into the pellet-die gap during pellet extrusion, leading to the formation of a lubrication layer that enables plug flow inside the pellet die.
Figure A.5: Longitudinal cross sections (x-ray tomograms obtained at the ANATOMIX beamline) of four pellets produced during the experimental trial at two different water contents after steam treatment, (a) two at low $\Delta\mathrm{mc}$ and (b) two at high $\Delta\mathrm{mc}$, 0.009 and 0.057 kg water per kg ingredients respectively. No indication for the flow direction is visible within the cross sections. The pellets flow as plugs through the pelleting die [60]. Scale bar indicates $0.5$ cm.
A.4 Friction, surface topology, and wear
In rheological experiments, it has been shown that various physical and chemical forces and constraints give rise to so-called slip-flow due to the displacement of the dispersed, solid-like, phase away from the moving boundary. This displacement of the dispersed phase, and hence the enrichment of the low-viscosity, the liquid-like, phase near the boundary, enhances material flow near the moving boundary [7]. In the case of stiff particles at the specified process conditions, particle size, excluded volume, and die-wall roughness are crucial properties of the system, determining whether lubrication layers form [7]. However, in the case of pelletizing soft and deformable particles of organic origin, parameters such as the excluded volume and mechanisms such as particle diffusion cannot be defined, as the confining pressure creates an essentially uniform density across the flowing zone. Particle size distributions only remain relevant for contact dynamics between the particles and the particle-wall interaction, as they set the intrinsic roughness scale at the paste-die interface.
Here, we extend the rheological observations to pellet extrusion, reasoning that the water layer between the flowing pellet and metal die, denoted as $\lambda$, reduces the direct contact area between the two sliding surfaces and thus reduces the apparent coefficient of friction. First, we present surface topology scans that show the addition of water during steam conditioning reduces wear on the pellet’s exterior surface, attributed to an increasing lubrication layer thickness between the sliding surfaces in the mixed lubrication regime, schematically represented in Fig. A.6. Secondly, we demonstrate how the addition of water through steam conditioning reduces the tribology, specifically the apparent coefficient of friction between the surfaces and contributes to lower mechanical energy use during pellet extrusion.
A.4.1 Surface topology and wear suggest mixed regime lubrication
Surface topology scans of pellets were conducted, to support the hypothesis of decreased contact area between the two sliding surfaces as a result of a water film with increasing thickness between the pellet and metal surface. Providing evidence for the type of lubrication regime during pellet extrusion. At low water contents, distinct striations are visible along the direction of extrusion (Fig. A.6b&d), indicating a boundary regime lubrication with direct contact between the pellet and die-wall. Under these conditions, the calculated lubrication layer thickness is relatively small in comparison to the expected roughness of the metal die inside the holes (see Fig 1d), resulting in surface wear due to protrusion of the metal asperities into the softer pellets’ surface. We were unable to obtain a measure for the surface roughness of the die during these experiments. This would require cutting of the stainless steel die into measurable pieces and thus destroy valuable equipment. However, the striations disappear when increasing the water content from $0.009$ kg water per kg ingredients during the conditioning process to $0.057$ kg water per kg ingredients (Fig. A.6c&e). This suggests that, in this case, the lubrication layer thickness increases, filling more cavities in the asperities between the pellet and the metal die, transitioning into the mixed lubrication regime [25]. In the mixed regime, the contact between the two surfaces is reduced, resulting in less wear on the pellet’s surface.
Figure A.6: Increasing lubrication layer thickness reduces the real contact area between the pellet and the die, reducing surface wear. a, Schematic representation of the pellet sliding along the metal surface of the pelleting die with asperities. Increasing lubrication layers are expected to increase the separation between the pellet and the die’s surface, reducing their contact area as expected in the mixed lubrication regime [25]. b & c, 3D surface topology scans of two pellets produced at low $\Delta\mathrm{mc}$ and high $\Delta\mathrm{mc}$, $0.009$ and $0.057$ kg water per kg ingredients respectively. d & e are the surface height profiles of the pellets scanned in (b) and (c). The black area’s correspond to peaks and valleys with a height or depth larger than $1.5\mathrm{\mu m}$. In (b) and (d) striations along the sliding direction are clearly visible, due to a thin lubrication layer. In (c) and (e) those striations are absent, due to filling of the cavities inside the pellet and the die, resulting in a reduced protrusion of the metal asperities into the extruded pellet.
A.4.2 Mixed lubrication and the apparent coefficient of friction
In the mixed lubrication regime, the apparent coefficient of friction, $\mu$ (-), is partially determined by the friction between the two solid surfaces, $\mu_{\mathrm{solid}}$ (-), and partially determined by the friction within the lubricating film, here water, $\mu_{\mathrm{water}}$ (-) [25]. The ratio between these two contributors changes with a change in the lubrication film thickness [26] (see Eq. 2):
$\mu=\mu_{\mathrm{solid}}X+\mu_{\mathrm{water}}(1-X)$ (2)
Here, $X$ (-), with values from 00 and $1$, represents the ratio between the load carried by the solid-solid contact interface with respect to the total contact load. The remainder, $(1-X)$, represents the load carried by the fluid film between the two surfaces. In the case of boundary lubrication, $X$ is equal to 1, and the apparent coefficient of friction, $\mu$, is fully determined by the solid-solid sliding friction. Non-lubricated pellet extrusion would occur during pellet production without the addition of water, steam, fat, oil, or other fluid lubricants. However, in the presence of a water film between the two sliding surfaces, $X$ reduces to values smaller than $1$, and the apparent coefficient of friction, determined from the experimental processing data, decreases.
A.4.3 Experimental coefficient of friction and energy use
According to the work by Holm et al., published in [45] and [47], a theoretical relationship exists that couples compression ratio - die length over die diameter, $L/D$ (-), and the coefficient of friction, $\mathrm{\mu}$, to the pressure required for extrusion during pelletization. In their works Holm et al. derive an exponential relationship (see eq. 3)
$P_{x}(x)=\frac{P_{0}}{\nu}\ \exp(4\mu\nu\frac{L}{D}-1)$ (3)
In which the extrusion pressure, $P_{x}$ $(\mathrm{Pa})$, is determined by a prestressing constant, $P_{0}$ $(\mathrm{Pa})$, the radial-to-axial stress ratio constant, or lateral pressure coefficient , $\nu$ (-), the coefficient of friction and the compression ratio of the die (see refs [45, 63] for further reference). By taking $\nu$ as a ingredient specific constant during pelletization, the pressure for extrusion increases exponentially with $\mu$ and $L/D$. Furthermore the extrusion pressure $(\mathrm{Nm^{-2}})$ and power consumption $(\mathrm{N\ ms^{-1}=W})$ are related through the volume flow rate $(\mathrm{m^{3}s^{-1}})$. Consequently, the power consumption of the pellet press, as well as the specific mechanical energy usage - the energy per unit mass - is expected to increase exponentially as a function of $\mu$ and $L/D$ as well (Fig. A.7).
Figure A.7: Increasing steam addition reduces the coefficient of friction between the pellet and the metal die. a, During pellet extrusion the specific mechanical energy usage is exponentially proportional to the coefficient of friction and compression ratio of the pelleting die. b Increasing the steam use during steam conditioning from a low moisture content ($0.035$ kg steam per kg ingredients) to a higher moisture content ($0.053$ kg steam per kg ingredients), results in a lower energy usage due to a decrease in the amount of friction. This is consistent with previous observations [37, 1].
Increasing the steam use during steam conditioning, achieved by controlling the temperature of the ingredient mixture after conditioning, leads to higher moisture content in the ingredient mixture and, as a result, reduces the mechanical energy demand during pellet extrusion [37, 1]. Our experimental data indicates that increasing steam use from $0.035$ kg steam per kg ingredients to $0.053$ kg steam per kg ingredients results in a reduction of the apparent coefficient of friction by approximately $16\%$ (see Fig. A.7). The thickness of the lubricating film is directly related to the increased separation of the two sliding surfaces (Fig 1d&e). By increasing the lubrication layer thickness the direct contact area between the metal die and the sliding pellet reduces, consequently $X$ reduces (see previous section) and the relative contribution of $\mu_{\mathrm{solid}}$ with respect to the apparent friction coefficient, $\mu$, decreases, reducing the extrusion sliding resistance. The contribution of the friction within the water film, is negligible (appendix C.2 - Energy dissipation in a liquid film). Hence our experimental observations supports the hypothesis that the increasing use of steam, and consequently the increase in moisture content, contributes to lowering the coefficient of friction between the metal die and the sliding pellet, through the formation of a lubrication layer in the mixed lubrication regime.
A.4.4 The effect of production rate on the mechanical energy use
In the mixed lubrication regime, the coefficient of friction decreases with increasing sliding velocity, though the rate of this decrease depends on specific factors such as surface hardness, roughness, lubricant properties, and load [64]. As discussed in previous sections, the amount of water added during the steam conditioning process determines whether pellet extrusion occurs in the boundary regime (with visible surface striations) or transitions into the mixed lubrication regime (where surface striations disappear). Although the operational space on the pellet press is limited compared to typical rheological devices, which can span several decades of shear rates, we were able to study the effect of production rate on mechanical energy use across a range of 100 to 500 kg/hr during pellet extrusion, with the addition of 0.057 kg of water per kg ingredients. As shown in Fig. A.8, a slight decrease in mechanical energy use is observed as production rate increases. While multiple factors may contribute to this trend, one possible explanation is the reduced coefficient of friction due to higher sliding speeds in the mixed lubrication regime, consistent with typical Stribeck behavior.
Figure A.8: Mechanical energy use during pellet extrusion decreases with production rate in the mixed lubrication regime. The transition into the mixed lubrication regime is marked by a reduction in mechanical energy use as the sliding velocity (or production rate) increases, corresponding to a decrease in the coefficient of friction. The error-bars correspond to the minimum and maximum values of two or three independent weighing samples taken during experimental runs.
A.5 Physical pellet quality
Figure A.9: Physical pellet quality can be optimized by tuning pellet temperature through steam and friction. a, During pelleting, energy is required for the production of steam (marked by “o”). In addition, electrical motor power is dissipated into the pellets (marked by “+”). Here, the total mechanical energy dissipated by the press, including the idle load of the pellet press (gross mechanical energy), is represented. The sum of the steam energy and mechanical energy determines the total energy used in pellet production (marked by “•”), which includes both conditioning and pelletizing. An optimum energy use exists, coupled with a normative pellet durability value (here $0.9$, equal to $90\%$), specifying the minimum amount of energy required to produce a high-quality pellet. This optimum is reached when pellet durability (b) no longer increases significantly with the invested energy during production (a). Below this optimum, pellets are still formed; however, the quality is insufficient for downstream processes, while above this optimum, the quality increases disproportionately with the energy investment (b). This pellet durability increase with increasing steam usage is consistent with [37].
Throughout pellet production, physical pellet quality is continually monitored for quality control. The key metric for physical pellet quality is pellet durability, which refers to the pellet’s ability to withstand handling and transportation without disintegrating [65]. High pellet durability indicates a strong resistance against breakage, which is critical in minimizing dust formation. Dust formation poses safety risks, including dust explosions, and health hazards associated with the inhalation of fine dust particles for both humans and animals. Therefore, achieving high durability is a crucial aspect of any pelleting process.
Producing a durable pellet with minimal energy use involves precise control over heat fluxes within the pellet production line. This control allows efficient adjustment of ingredient and pellet temperatures to facilitate bond formation and produce durable pellets (see Fig. A.9b). The heating process begins with the addition of steam during steam conditioning, followed by heating through friction during pellet extrusion. Increasing steam usage enables a reduction in mechanical energy (see Fig. 3b and Fig. A.9a). However, it is essential to note that steam production itself is energy-intensive (see Fig. A.9a), here estimated through assuming a $100\%$ efficiency of the steam-boiler. Therefore, finding a balance between steam and mechanical energy fluxes is crucial for optimizing pellet production for energy efficiency. Beyond a certain temperature threshold for a specified ingredient composition, further increases in total energy input (see Fig. A.9a - ”Total energy use”) yield no significant durability improvements (see Fig. A.9b). Above this optimum process temperature, production costs increase significantly with continued steam usage. To mitigate these costs, replacing part of the added steam with tap water during conditioning might be considered. Tap water can contribute to lubrication, reducing mechanical energy use without the energy costs associated with an increased steam production. Therefore active monitoring of the lubrication layer, the mechanical energy use and product temperatures, is a requirement for future process optimization during pellet extrusion.
Appendix B Methods
B.1 Materials
A total of 18,000 kg of milled material was obtained from Research Diet Services (Wijk bij Duurstede, The Netherlands). The material was ground using a hammer mill equipped with 4-mm screens. The material, a 50/50% w/w mixture of cleaned maize kernels and low-sugar sugar beet pulp, was delivered in two separate orders: the first batch (Batch 1) contained 7,000 kg, and the second batch (Batch 2) contained 11,000 kg. Batch 1 was used to obtain the results presented in all figures except Fig. A.7, Batch 2 was used solely for the results presented in Fig. A.7. The products, being of organic/natural origin, may differ in their chemical compositional between Batch 1 and Batch 2. However, these variations were not the focus of this study, therefore the chemical composition was not analyzed.
B.2 Pellet production and process monitoring
A RMP200 ring-die pelletizer (Münch Edelstahl GmbH, Hilden Germany) served as the pilot-scale pellet extrusion system. Configured with two rollers, the ring-die featured 6 mm diameter die-holes and a compression ratio (L/D) of 12, corresponding to a channel length of 72 mm. In addition to the 6 mm L/D 12 die, three additional dies with 6 mm diameter holes were used for the results presented in Fig. A.7, with L/D ratio’s 6, 9.3 and 16 respectively. The roller-die gap was set at 0.20 mm, and the die rotation frequency maintained at 5 Hz.
Steam injection and mixing, facilitated by a single-screw steam conditioner (Münch Edelstahl GmbH, Hilden Germany), adjusted the water content of the ingredient mixture prior to pelletization. Super-heated steam was injected at 140°C and used to set the ingredient temperature inside the steam conditioner.
Each run was initialized with a 20 minutes period, during which feeder speed and steam injection rate were adjusted to obtain steady-state. This steady-state was confirmed by monitoring of the electrical current use (A, ampere) of the pellet press using a Janitza Power Analyzer UMG-604 (Janitza, Lahnau, Germany).
Registered process parameters include ingredient mixture temperature and moisture content before and after steam conditioning, pellet temperature and moisture content after pellet extrusion, electrical current, and production capacity (kg/h). From which the change in moisture content ($\Delta\mathrm{mc}$, in kg water per kg ingredients) during the steam conditioning process, the change in temperature over the extrusion die ($\Delta\mathrm{T_{die}}$) and the specific mechanical energy (SME) were calculated. The temperature measurements were performed as described in [44], the moisture content was determined as described in [54] and the production capacity (also: mass-flow-rate) was determined as described in [44], using a 2 minute measurement period.
The gross and net SME (kWh/t or J/kg) were determined according to the method described in [44], where the gross SME is calculated using the full-load current and the net SME is calculated by subtracting the idle-load current from the the full-load current. This enables us to differentiate between the total amount of energy dissipated (idle-load plus pellet extrusion) by the press and the amount of energy dissipated in the pellet extrusion process only.
During the processing of Batch 2, additionally, the steam-flux (kg steam per kg ingredients) was recorded (H250, Krohne, Duisburg, Germany), instead of the change in moisture content before and after conditioning.
During the processing of Batch 1, the effect of steam dosing on pellet extrusion was evaluated by conducting five steady-state runs in random order at constant production capacity (approx. 250 kg/h) and variable steam injection rate, thereby tuning $\Delta\mathrm{mc}$ as specified in Fig. 1d. During the processing of Batch 2, the effect of L/D ratio and steam dosage on the specific mechanical energy use was evaluated using a constant production capacity (approx. 250 kg/h) and at two levels of $\Delta\mathrm{mc}$ (0.035 and 0.053 kg steam per kg ingredients, respectively), while changing the die-configuration as specified in Fig. A.7b.
B.3 Drying setup and analysis
The lubrication layer thickness was determined according to the method described in [34], using a model drying setup. The drying setup, consists of a 3D-printed air-tunnel, into which a controlled air-flow (40 L/min, FL-2044, Omega Engineering Limited, Manchester, United Kingdom) is supplied from one side. The air passes through a 3D printed (PolyLite ASA - black, Polymaker) sample container containing a sample of pellets, which is placed on top of an analytical balance (ME204, Mettler-Toledo B.V., Tiel, the Netherlands). During the drying process, the temperature and relative humidity (RH) of the air entering and exiting the sample container with pellets over time was recorded using a TSP01 RevB sensor (Thorlabs Inc.).
The sample container (6x6x5 cm) was equipped with two sheets of stainless steel woven wire mesh (6x6 cm) with aperture of 2.4 mm, positioned at opposite sides of the container, allowing the air to pass through the filled sample container, whilst retaining the pellets. The input airline was connected to a compressed air line, available in the process hall. The flow rate of 40 L/min, corresponds to an air velocity of 0.19 m/s upon entering the sample container. Samples of approximately 100g of fresh pellets were collected in the sample container directly below the pellet press. The mass of the sample container was recorded for a period of 15 minutes at a sampling frequency of 1 Hz, after which the moisture ratio ($\mathrm{MR}$, see Eq. 6) was determined. The air dried material was transferred to a plastic bag and sealed immediately. The moisture content (kg water per kg pellet) of the pellets after drying was subsequently determined according to the method described in [54].
Model drying experiments were performed three times (20 minutes apart) during the steady-state of the pelleting press by placement container with freshly sampled pellets in the drying setup. The saturation time period, $t_{\mathrm{sat}}$, was determined visually from the RH-curves, and is defined as the time period during which the RH-value is larger than 95%, indicated by the green area in Fig. 1b. The amount of evaporated water (kg water per kg pellet) was determined through multiplication of the averaged moisture ratio ($\mathrm{MR}$, unitless) over time and the averaged change in moisture content (kg water per kg pellet). The latter is determined as the difference in the moisture content at the start ($\mathrm{mc}$ right after pellet extrusion) and at the end of the drying experiment ($\mathrm{mc}$ of the pellets collected, bagged and sealed).
B.4 Calculating $\lambda$
The lubrication layer thickness is calculated from the amount of evaporated water at time $t_{\mathrm{sat}}$ during the drying experiment [34]. The amount of evaporated water ($X$, kg water per kg pellet) is converted to an annular volume of water ($V_{\mathrm{water}}$) on top of the pellets surface, with $V_{\mathrm{water}}=\pi l(R^{2}-r^{2})$. The lubrication layer thickness $\lambda$ is calculated using $\lambda=R-r$, in which $R$ is the outer radius of the annulus (water layer plus pellet radius) and $r$ the radius of a pellet. $R$ is unknown but is determined from $R^{2}=(V_{\mathrm{water}}/V_{\mathrm{pellet}}+1)r^{2}$ using $V_{\mathrm{water}}/V_{\mathrm{pellet}}=X/\rho_{\mathrm{water}}\times\rho_{\mathrm{pellet}}$. The latter equation converts the amount of evaporated water in kg water per kg pellet, into the evaporated water volume per volume pellet. To calculate $R$, the density of the pellet, $\rho_{\mathrm{pellet}}$ is assumed to be $1200\ \mathrm{kg_{\mathrm{pellet}}\ m^{-3}}$[66], the radius of the a pellet is approximated by the inner radius of the extrusion channel ($D/2$) here $3\ \mathrm{mm}$ (see Fig. A.2), and the density of water $\rho_{\mathrm{water}}$ is approximated at $1000\ \mathrm{kg_{\mathrm{water}}\ m^{-3}}$.
B.5 Rehydration of pellets
As a control drying experiments, a sample of dried pellets were rehydrated through hygroscopic vapor sorption. This sample of approximately 300 grams of pellets was stored in a container maintained at >95% RH for over 1 week, before being dried in the model dryer, using the same drying conditions as during the drying of fresh pellets. Two drying runs were performed, one lasting 3.5 hours and the other approximately 15 hours (yellow lines in Fig. A.1c).
B.6 Physical pellet quality
Pellet durability was determined using a Holmen NHP100 (Tekpro, USA) tester with a 3-mm screen. During each steady-state run, three pellet samples of approximately 2 kg were sampled directly after leaving the press. The sampling was performed approximately 20 minutes apart. The 2 kg samples were subsequently cooled at ambient air for 10 min and transferred into two plastic pots (1 L). Approximately 100 grams of pellets, from each pot, underwent testing in the rotating drum mechanism, simulating mechanical stress [65]. The Pellet Durability Index (PDI) was calculated based on weight loss after tumbling, indicating the percentage of intact pellets after testing. A high pellet durability (%), indicates a high resistivity of pellets against abrasion during handling and transport.
B.7 Synchrotron-based X-ray microtomography
Computed tomography (CT) scans were conducted on pellets produced under various process conditions utilizing the ANATOMIX beamline at the synchrotron SOLEIL. Each pellet was securely mounted in custom 3D printed sample holders, allowing for individual scanning. These holders were positioned in the Huber goniometer head, maintaining a consistent distance of 20 cm from the scintillator. Image acquisition utilized a 2048x2048 pixel detector with a pixel size of 3.07 µm, such that the pellet (approximately 6 mm in diameter) fits entirely within the detectors field of view. The scanning process involved a rotation step of 0.09°, resulting in 2000 projections collected over a 180° rotation. Reconstruction of the acquired data into a 3D image volume was performed using a Paganin filter. In Fig. A.5, four scans are shown of four individual pellets produced at two steam injection levels (2x2).
B.8 Density profile
The radial density profile of an X-ray tomogram was determined using a custom MATLAB (R2023a) code. The 3D image volume of a pellet was segmented into 30 equidistant slices along the axial direction (top to bottom), with each slice processed individually.
Initially, the center of the pellet was identified, and subsequently, an Euclidean distance map was generated. This map enabled the calculation of the distance of each pixel in the slice relative to the pellet’s center point, measured in pixels. Then, the intensity values of each pixel at a given distance (1-pixel bins) were averaged. This process was repeated for all 30 slices.
To ensure data quality, the average intensity of each image slice was utilized as a threshold to discard slices with low X-ray photon counts, resulting from drift in the alignment of the monochromatic beam with respect to the sample and the detector. The remaining slices were then averaged to obtain the final average density profile along the radial coordinate, representing the distance from the pellet’s center point.
B.9 Scanning electron microscopy
SEM was performed on a FEI Magellan 400 (FEI Electron Optics B.V., The Netherlands) using an acceleration voltage of 2.0 kV. The pellets were mounted on aluminum stubs with double-sided adhesive, conductive carbon tape. Pellets were cut along the longitudinal axis. The outer and inner pellet surface was oriented upward and sputter-coated with 12 nm of Tungsten under a 45° angle (Leica EM SCD500, Leica Microsystems, The Netherlands).
B.10 Surface characterisation
An automatic, noncontact laser confocal microscope (VK-X1000 Series; KEYENCE, Japan) was used to analyze surface topology of pellets at ×20 magnification with a corresponding pixel size of 0.686 $\mathrm{\mu m}$. The surface height profile was exported in the KEYENCE software and further analyzed using a custom code in MATLAB (R2020b). In MATLAB the height data was treated as image data (768x1024 pixels). A 3D surface profile was generated by converting the height data into a stereolithography (.STL) file. The background curvature, the cylinder shape of the pellets, was determined and afterwards subtracted from the original height data by using a Gaussian filter with a large value for sigma. The value for sigma (= 60) was chosen to smooth the height data over a large area, which prevents smoothing the much smaller features at the pellets surface. The flattened height data was then trimmed by 60 pixels along all four edges to remove artifacts at the image boundaries (648x904 pixels). The trimmed and flattened height data was used for visualization of the height profiles (see Fig. A.6).
Appendix C appendix
C.1 Simulation details COMSOL
Numerical simulations were performed using COMSOL Multiphysics software to analyze the drying behavior of pellets. This study focused on the impact of the redistribution of water along the radial coordinate due to radial phase migration. The objective of these simulations was to qualitatively compare the experimental drying data with the potential impact, as revealed through numerical modeling, of radial phase migration on drying behavior. It is essential to note that the experimental data were obtained through the drying of a container with approximately 100 grams of sample material, whereas the numerical effect of radial redistribution of moisture was simulated on a single pellet level. Evaporation from the top and bottom ends of a pellet was ignored. Therefore, drying of a pellet was simulated along a 1D line of length $r_{p}$ ($3\times 10^{-3}\ \mathrm{m}$), the radius of a pellet produced in a 6 mm pellet die, with a symmetry axis at $r=0$.
The governing partial differential equation for the transport of moisture, $M$, throughout the pellet was based on Fick’s law:
$\frac{\partial M}{\partial t}=\frac{1}{r}\frac{\partial}{\partial r}\left(rD_{\rm eff}\frac{\partial M}{\partial r}\right)$ (4)
where the rate of moisture loss $\partial M/\partial t$ is controlled by the moisture diffusivity inside the pellet, $D_{\rm eff}$.
To solve eq. 4, a convective boundary condition was used, formulated as an analogue to the Hertz-Knudsel-Schrage equation, wherein the gradient in vapor pressure drives the evaporation process, as expressed by equation 5:
$-D_{\rm eff}\left.\frac{\partial M}{\partial r}\right|_{r=r_{p}}=k_{c}\left(0-M\right)$ (5)
where $k_{c}$ is the mass transfer coefficient. The concentration outside of the pellet was set to 0 based on the following rationale: During pellet production, water is added to the ingredient mixture through the condensation of steam during steam conditioning. This added water elevates the moisture content within the pellets above its equilibrium moisture content ($M\geq M_{\mathrm{eq}}$), the moisture content of the ingredient mixture pre-conditioning. The added water will evaporate until the equilibrium moisture content is restored ($M=M_{\mathrm{eq}}$). This equilibrium state implies that there is no longer a gradient in partial vapor pressure between the air and the ingredient mixture, resulting in equal rates of evaporation and condensation. Consequently, the effective external moisture content was set to 0 (as per equation 5), indicating that all added water will eventually evaporate as the sign of $M$ is positive (reflecting the addition of water during pellet production).
All simulations were initialized with an initial moisture content of $+2.5\%$ (kg water per kg material), a typical increase in moisture content during steam conditioning.
The value for $D_{\rm eff}$ was estimated at $5\times 10^{-11}\ \mathrm{m^{2}\ s^{-1}}$ which matches typical values reported in literature for foodstuffs [56] and closely matches the experimental drying rate of the uniformly rehydrated pellets shown by the yellow lines in Fig. A.1. The value of $k_{c}\ (2\times 10^{-2}\ \mathrm{m\ s^{-1})}$ was selected such that the mass Biot number $\mathrm{Bi_{m}}=\frac{k_{c}\times r_{p}}{D_{\rm eff}}=1.2\times 10^{6}\gg 1$. Therefore, the rate of diffusion within the pellet is the evaporation rate-limiting factor, as can be expected for hygroscopic organic material that slowly absorbs or releases water depending on the relative humidity of the air in the environment [54].
[…]
$\mathrm{MR}=\frac{X(t)-X(t=900\mathrm{s})}{X(t=0\mathrm{s})-X(t=900\mathrm{s})}$ (6)
C.2 Energy dissipation in a liquid film
The surface topology scans, exhibiting clear striation patterns upon the addition of minimal water during pellet manufacturing and the disappearance thereof with increased water addition, suggest that we are within the mixed lubrication regime. In this regime, where $X$ ranges between $1$ and 00, the work of friction is predominantly determined by solid-solid contact friction (see eq. 2). The exact contributions of $\mu_{\mathrm{solid}}$ and $\mu_{\mathrm{water}}$ at operational processes are unknown. However, the contribution of friction in the water film can be estimated by assuming a full separation of the two sliding surfaces, in which case all friction is assumed to result from shearing of the lubrication layer. Through careful estimates, we can determine the specific energy dissipation due to the shearing of the water film (see eq. 7):
$\frac{P_{\mathrm{diss}}}{Q_{\mathrm{m}}}=\frac{V_{\mathrm{water}}\eta_{\mathrm{water}}\dot{\gamma}_{\mathrm{water}}^{2}}{\dot{V}_{\mathrm{pellet}}\rho_{\mathrm{pellet}}}$ (7)
Here, $P_{\mathrm{diss}}$ is the energy dissipation in the water film, $Q_{\mathrm{m}}$ is the mass-flow-rate of the pellet through the die-hole. $V_{\mathrm{water}}$ is the volume of the sheared water, an annulus with outer radius $r$ (here $3\ \mathrm{mm}$) and inner radius $r-\lambda$, using the height of the water film, $\lambda$ (here approximated by $10\ \mathrm{\mu m}$) and die-hole channel length $L$ (here $72\ \mathrm{mm}$). The viscosity of water, $\eta_{\mathrm{water}}$, at room temperature is approximately $1\ \mathrm{mPa\ s}$. The shear rate $\dot{\gamma}_{\mathrm{water}}$ is calculated from the extrusion velocity, $v$ (approximately $1\ \mathrm{ms^{-1}}$), the velocity at which the pellet moves through the die when the roller passes over the die-hole, and the height of the water film, $\lambda$. Thus, the shear rate is approximately $10^{5}\ \mathrm{s^{-1}}$. $\dot{V}_{\mathrm{pellet}}$ is the volumetric flow rate of a cylindrical pellet through the die-hole, for a pellet with radius $r-\lambda$ and using a displacement rate equal to the extrusion velocity ($\dot{V}_{\mathrm{pellet}}=\pi(r-\lambda)^{2}v$). And the pellet’s density, $\rho_{\mathrm{pellet}}$ (here approximated at $10^{3}\ \mathrm{kg\ m^{-3}}$).
The specific energy dissipation, in the case where all the shear stresses would act on a water-based lubrication layer of $10\ \mathrm{\mu m}$ thickness, is estimated to be $5\ \mathrm{J\ kg^{-1}}$. However, the experimental data shows an net energy dissipation of more than $60\ \mathrm{kJ\ kg^{-1}}$ upon the presence of a $10\ \mathrm{\mu m}$ lubrication layer during the pelleting trial (see Fig. 3). Therefore, we conclude that we are in the mixed lubrication regime, in which the predominant part of the apparent friction is determined by the solid-solid sliding friction, $\mu_{\mathrm{solid}}$, and not by the friction within the water film, $\mu_{\mathrm{water}}$, see eq. 2.